Based on a​ poll, among adults who regret getting​ tattoos, 17​% say that they were too young when they got their tattoos. Assume that nine adults who regret getting tattoos are randomly​ selected, and find the indicated probability.



1
Expert's answer
2021-12-14T13:14:58-0500

Based on a poll, among adults who regret getting tattoos, 17 % say that they were too young when they got their tattoos. Assume that nine adults who regret getting tattoos are randomly selected, and find the indicated probability. Find the probability that none of the selected adults say that they were too young to get tattoos

p=17%=0.17

n=9

x=0

Binomial distribution

"P(X=x) = C^{9}_x(0.17)^x(1-0.17)^{9-x} \\

P(X=0) = C^{9}_0(0.17)^0(1-0.17)^{9-0} \\

= 0.83^9 \\

= 0.1869"